OpenStudy (anonymous):

A bag contains 3 blue marbles, 9 green marbles, and 11 green marbles. Twice you draw a marble and place it. Find P(blue, then green) A.27/529 B.27/23 C.15/529 D.12/23

OpenStudy (y1066380):

Total number of marbles=3+9+11=23 P(blue, then green) =[get blue in the first draw][get green in the second draw] =3/23*9/22 =27/506 Am I correct?